3. Carol makes $1,900 per month and has an opportunity to save $100 per month at an APR of 3.25% in a 401K plan through work. She plans to retire in 40 years. a. If Carol’s company offered a 50% matching, use Excel to determine how much her investment will be worth in 40 years. i. Write the Excel formula that gives its future value _________________________________________________________ ii. Find the balance in the account after 40 years: _________________________ b. How much will Carol have deposited in the 401K over the 40 years? c. How much would the company contribute to Carol’s 401K over the 40 years? d. At retirement how much more is in the account than Carol deposited? e. Carol is not sure that she can afford to invest in the 401K. What would you say to convince her that she is making a good financial decision?
Answer #1

Answer:

Given that

a) Excel future value= FV (rate = 3.25% /12 n per = 40 x 12 pmt = 100 +50% x 100 pv = 0,0)

= $ 147, 481.17

b) carol company contribution = 100 x 50% x 12 x 40 = $ 24000

c) carol deposited = 24000 x 2 = 48000

Difference = 147481.17 - 48000 = 99,481.17

--------------------

If you have any query regarding the answer please as me in the comment I am here for help you. Please do not direct thumbs down just ask if you have any query. And if you like my work then please appreciates with up vote
Thank you Very much.